OpenStudy (anonymous):

expand sin3(theta) in terms of cos and sin

OpenStudy (anonymous):

sin3x=3sinx-4(sinx)^3

OpenStudy (anonymous):

take sinx common =sinx(3-4(sinx)^2) =sinx(3-3(sinx)^2-(sinx)^2) =sinx(3(1-(sinx)^2)-(sinx)^2) =sinx(3*(cosx)^2-(sinx)^2) solve further now